# Nearly Tight Black-Box Auditing of Differentially Private Machine Learning

### v1

@@ 5.1 Experimental Setup
Datasets. We experiment with the MNIST [22] and CIFAR-10 [21] datasets. The former includes 60,000 training and 10,000 testing 28x28 grayscale images in one of ten classes (hand-written digits), Often referred to as a “toy dataset,” is is commonly used to benchmark DP machine learning models. CIFAR-10 is a more complex dataset containing 50,000 training and 10,000 testing 32x32 RGB images, also consisting of ten classes. Note that the accuracy of DP models trained on CIFAR-10 has only recently started to approach non-private model performance [10, 33]. Whether on MNIST or CIFAR-10, prior black-box audits (using average-case initial parameters) have not been tight. More precisely, for theoretical $\varepsilon=8.0$, De et al. [10] and Nasr et al. [26] only achieve empirical estimates of $\varepsilon_{emp}\approx 2$ and $\varepsilon_{emp}\approx 1.6$, respectively, on MNIST and CIFAR-10.
To ensure a fair comparison between the average-case and worst-case initial parameter settings, we split the training data in two and privately train on only half of each dataset (30,000 images for MNIST and 25,000 for CIFAR-10). The other half of the training dataset is used as the auxiliary dataset to non-privately pre-train the models and craft the worst-case initial parameters (see Section 4). We note that for CIFAR-10, we additionally augment the pre-training dataset with the ImageNet-32 dataset to boost the model accuracy further (see Appendix A for further details).
Models. When training the model fully, we use a (shallow) Convolutional Neural Network (CNN) for both datasets. We do so as we find accuracy to be poor with models like Logistic Regression and Fully Connected Neural Networks, and training hundreds of deep neural networks is impractical due to large computational cost. Therefore, we draw on prior work by Dörmann et al. [13] and train shallow CNNs that achieve acceptable model accuracy while being relatively fast to train. (We report exact model architectures in Appendix A). Whereas, for the last layer-only fine-tuning setting, we follow [10] and use the Wide-ResNet (WRN-28-10) [37] model pre-trained on ImageNet-32 [11].
Experimental Testbed. All our experiments are run on a cluster using 4 NVIDIA A100 GPUs, 1 CPU core, and 12GB of RAM. Auditing just one model took approximately 16 GPU hours on MNIST and 50 GPU hours on CIFAR-10.
@@ 5.2 Full Model Training
Model Accuracy. In Table 1, we report the accuracy of the final models for the different values of $\varepsilon$-s we experiment with. While accuracy on MNIST is very good ($\geq 95\%$ for all $\varepsilon$), for CIFAR-10, we are relatively far from the state of the art (${\approx}50\%$ at $\varepsilon=10.0$ vs. ${\approx}80\%$ at $\varepsilon=8.0$ in [10]). This is due to two main reasons. First, as we need to train hundreds of models, we are computationally limited to using shallow CNNs rather than deep neural networks. Second, we need to consider a much larger batch size to ease auditing [26]; thus, the models converge much more slowly. In Appendix B, we show that, given enough iterations ($T=1{,}000$), the model reaches an accuracy of ${\approx}70\%$ at $\varepsilon=10.0$, almost approaching state-of-the-art. As a result, we audit models striking a reasonable trade-off between utility and computational efficiency, as discussed above.
Impact of initialization. In Figure 1, we compare the difference in $\varepsilon_{emp}$ between worst-case initial parameters and average-case parameters (as done in prior work [10, 18, 26]). In the latter, all models are initialized to the same, randomly chosen $\theta_{0}$, while, in the former, to a $\theta_{0}$ from a model pre-trained on an auxiliary dataset. In the rest of this section, the target sample is set as the blank sample as done in previous work [10, 26].
With worst-case initial parameters, we achieve significantly tighter audits for both MNIST and CIFAR-10. Specifically, at $\varepsilon=1.0,2.0,4.0,10.0$, the empirical privacy leakage estimates reach $\varepsilon_{emp}=0.08,0.54,1.89,6.48$ and $\varepsilon_{emp}=0.00,0.18,1.43,4.96$ for MNIST and CIFAR-10, respectively. In comparison, with average-case initial parameters, the empirical privacy leakage estimates were significantly looser, reaching only $\varepsilon_{emp}=0.00,0.06,0.61,3.41$ and $\varepsilon_{emp}=0.07,0.02,0.20,0.69$ for MNIST and CIFAR-10, respectively.
@@
Impact of size of dataset. In the auditing literature, empirical privacy leakage is often commonly estimated on (smaller) dataset samples both to achieve tighter audits and for computational efficiency reasons [10]. In Figure 3, we evaluate the impact of the dataset size ($n$) on the tightness of auditing. While smaller datasets generally yield tighter audits, the size’s impact also depends on the dataset itself. For instance, at $\varepsilon=10.0$, auditing with $n=100$ results in $\varepsilon_{emp}=7.41$ and $7.01$, for MNIST and CIFAR-10, respectively, compared to $\varepsilon_{emp}=6.48$ and $4.96$ on the full dataset. In other words, with MNIST, the dataset size does not significantly affect tightness. However, for CIFAR-10, smaller datasets significantly improve the audit’s tightness significantly.
Figure 5: Auditing models when fine-tuning only the last layer using average-case and worst-case initialization of models.
Sensitivity to clipping norm. Finally, we evaluate the impact of the hyper-parameters on the tightness of black-box audits. In the black-box setting, unlike in white-box, the adversary cannot arbitrarily insert gradients that are of the same scale as the gradient clipping norm, $C$ [26]. Thus, they are restricted to gradients that naturally arise from samples. In Figure 4, we report the $\varepsilon_{emp}$ values obtained with varying clipping norms, $C$. For computational efficiency, the results on CIFAR-10 are for models trained on datasets of size $n=1{,}000$.
@@ 5.3 Fine-Tuning the Last Layer Only
Once again, we compare the tightness of the audits with respect to the initial model parameters. In Figure 5, we report the empirical privacy leakage estimates obtained at various levels of $\varepsilon$ when the last layer is initialized to average- and worst-case (through pre-training) initial model parameters. Similar to the full model training setting, model initialization can impact audit tightness, albeit not as significantly. Specifically, at $\varepsilon=10.0$, we obtain $\varepsilon_{emp}=7.69$ and $8.30$, respectively, when the last layers are initialized to average- and worst-case initial model parameters.
Overall, the audits in this setting are considerably tight, and increasingly so for larger $\varepsilon$, as we obtain empirical privacy leakage estimates of $\varepsilon_{emp}=0.15,0.99,2.65,8.30$ for $\varepsilon=1.0,2.0,4.0,10.0$, respectively. The maximum $\varepsilon_{emp}$ across five independent runs is also much tighter, namely, $\varepsilon_{emp}=0.69,1.75,3.90,10.10$ for $\varepsilon=1.0,2.0,4.0,10.0$, respectively. Although maximum $\varepsilon_{emp}=10.10$ at $\varepsilon=10.0$ suggests a DP violation, this is, in fact, within the standard deviation of ±$\pm 1.00$.
@@ 6 Conclusion
This paper presented a novel auditing procedure for Differentially Private Stochastic Gradient Descent (DP-SGD). By crafting worst-case initial model parameters, we achieved nearly tight empirical privacy leakage estimates for DP-SGD in the black-box model and for natural (not adversarially crafted) datasets. At $\varepsilon=10.0$, we achieve nearly tight estimates of $\varepsilon_{emp}=7.21$ and $6.95$ for datasets consisting of 1,000 samples from MNIST and CIFAR-10, respectively. While we achieve slightly weaker estimates of $\varepsilon_{emp}=6.48$ and $4.96$ on the full datasets, these still represent a roughly 3x improvement over the estimates achieved in prior work [10, 26].
@@
Relatedly, recent work [3, 32] has begun to focus on auditing within a single training run, although it has thus far achieved only loose empirical estimates. Finally, in future work, we plan to investigate whether the audits can be tight in the black-box model even when the batch size $B<n$.
@@
Appendix A Model Architectures and
Hyper-parameters
@@ Appendix A Model Architectures and
Hyper-parameter tuning. To achieve the best model utilities, we tune the hyper-parameters of DP-SGD (i.e., number of iterations, $T$ and learning rate, $\eta$). For MNIST, we train for $T=100$ iterations, with a learning rate of $\eta=4$ for both average-case and worst-case initial model parameters. For CIFAR-10, we train for $T=200$ iterations, with a learning rate of $\eta=2$ for the average-case initial model parameter setting. For the worst-case initial model parameter setting, we use a learning rate of $\eta=1$ instead, as this resulted in the best model performance for CIFAR-10. When fine-tuning just the last layer, we train for $T=25$ iterations, with a learning rate of $10.0$. Note that, for the experiments varying the dataset size, the learning rate was scaled accordingly so that the size of the dataset does not affect the “influence” of each sample (i.e., $\frac{\eta}{B}$). All clipping norms are set to $C=1.0$ as done in prior work [10], and batch sizes are set to the dataset size $B=n$ to ease auditing [26]. Finally, the noise multiplier $\sigma$ is calculated from the batch size and number of iterations using the Privacy loss Random Variable accountant provided by Opacus [36].
Worst-case initial model parameters, $\theta_{0}$. To craft the worst-case initial model parameters for MNIST, we pre-train the model on half of the full dataset (reserving the other half to be used for DP-SGD) for 5 epochs with a batch size of 32 and a learning rate of 0.01. For CIFAR-10, we first pre-train the model on the ImageNet-32 dataset for 300 epochs with batch size 128 and learning rate 0.1. Then, we (non-privately) fine-tune the model on half of the full dataset for 100 epochs, with a batch size of 256 and a learning rate of 0.1.

### v3

@@ 5.1 Experimental Setup
Datasets.
We experiment with the MNIST [26] and CIFAR-10 [25] datasets. The former includes 60,000 training and 10,000 testing 28x28 grayscale images in one of ten classes (hand-written digits), Often referred to as a “toy dataset,” is commonly used to benchmark DP machine learning models. CIFAR-10 is a more complex dataset containing 50,000 training and 10,000 testing 32x32 RGB images, also consisting of ten classes. Note that the accuracy of DP models trained on CIFAR-10 has only recently started to approach non-private model performance [13, 38]. Whether on MNIST or CIFAR-10, prior black-box audits (using average-case initial parameters) have not been tight. More precisely, for theoretical $\varepsilon=8.0$, De et al. [13] and Nasr et al. [31] only achieve empirical estimates of $\varepsilon_{emp}\approx 2$ and $\varepsilon_{emp}\approx 1.6$, respectively, on MNIST and CIFAR-10.
To ensure a fair comparison between the average-case and worst-case initial parameter settings, we split the training data in two and privately train on only half of each dataset (30,000 images for MNIST and 25,000 for CIFAR-10). The other half of the training dataset is used as the auxiliary dataset to non-privately pre-train the models and craft the worst-case initial parameters (see Section 4). We note that for CIFAR-10, we additionally augment the pre-training dataset with the CIFAR-100 dataset to boost the model accuracy further (see Appendix A for further details).
Models.
When training the model fully, we use a (shallow) Convolutional Neural Network (CNN) for both datasets. We do so as we find accuracy to be poor with models like Logistic Regression and Fully Connected Neural Networks, and training hundreds of deep neural networks is impractical due to large computational cost. Therefore, we draw on prior work by Dörmann et al. [16] and train shallow CNNs that achieve acceptable model accuracy while being relatively fast to train. (We report exact model architectures in Appendix A). Whereas, for the last layer-only fine-tuning setting, we follow [13] and use the Wide-ResNet (WRN-28-10) [44] model pre-trained on ImageNet-32 [14].
Experimental Testbed.
All our experiments are run on a cluster using 4 NVIDIA A100 GPUs, 64 CPU cores, and 100GB of RAM. Auditing just one model took approximately 16 GPU hours on MNIST and 50 GPU hours on CIFAR-10.
@@ 5.2 Full Model Training
Model Accuracy.
In Table 1, we report the accuracy of the final models for the different values of $\varepsilon$-s we experiment with. While accuracy on MNIST is very good ($\geq 95\%$ for all $\varepsilon$), for CIFAR-10, we are relatively far from the state of the art (${\approx}50\%$ at $\varepsilon=10.0$ vs. ${\approx}80\%$ at $\varepsilon=8.0$ in [13]). This is due to two main reasons. First, as we need to train hundreds of models, we are computationally limited to using shallow CNNs rather than deep neural networks. Second, we need to consider a much larger batch size to ease auditing [31]; thus, the models converge much more slowly. In Appendix B, we show that, given enough iterations ($T=1{,}000$), the model reaches an accuracy of ${\approx}70\%$ at $\varepsilon=10.0$, almost approaching state-of-the-art. As a result, we audit models striking a reasonable trade-off between utility and computational efficiency, as discussed above.
Impact of initialization.
In Figure 1, we compare the difference in $\varepsilon_{emp}$ between worst-case initial parameters and average-case parameters (as done in prior work [13, 22, 31]). In the latter, all models are initialized to the same, randomly chosen $\theta_{0}$ (using Xavier initialization [20]), while, in the former, to a $\theta_{0}$ from a model pre-trained on an auxiliary dataset. In the rest of this section, the target sample is set as the blank sample, which is known to produce the tightest audits in previous work [13, 31].
With worst-case initial parameters, we achieve significantly tighter audits for both MNIST and CIFAR-10. Specifically, at $\varepsilon=1.0,2.0,4.0,10.0$, the empirical privacy leakage estimates reach $\varepsilon_{emp}=0.08,0.54,1.89,6.48$ and $\varepsilon_{emp}=0.00,0.18,1.43,4.96$ for MNIST and CIFAR-10, respectively. In comparison, with average-case initial parameters as in prior work, the empirical privacy leakage estimates were significantly looser, reaching only $\varepsilon_{emp}=0.00,0.06,0.61,3.41$ and $\varepsilon_{emp}=0.07,0.02,0.20,0.69$ for MNIST and CIFAR-10, respectively. Notably, at larger privacy levels $\varepsilon=4.0$ and $\varepsilon=10.0$, for both MNIST and CIFAR-10, the empirical privacy leakage estimates when auditing with worst-case initial parameters are not only larger than the average-case, but are outside of the $\pm$ s.d. range. This shows that our auditing procedure not only produces significantly tighter audits than prior work, especially at larger $\varepsilon$s, but also that, at lower $\varepsilon$-s, our evaluations are currently limited by the number of models used to audit (resulting in large s.d.).
@@
Figure 5: Auditing models when fine-tuning only the last layer using average-case and worst-case initialization of models at various levels of $\varepsilon$.
Impact of size of dataset.
In the auditing literature, empirical privacy leakage is often commonly estimated on (smaller) dataset samples both to achieve tighter audits and for computational efficiency reasons [13]. In Figure 3, we evaluate the impact of the dataset size ($n$) on the tightness of auditing. While smaller datasets generally yield tighter audits, the size’s impact also depends on the dataset itself. For instance, at $\varepsilon=10.0$, auditing with $n=100$ results in $\varepsilon_{emp}=7.41$ and $7.01$, for MNIST and CIFAR-10, respectively, compared to $\varepsilon_{emp}=6.48$ and $4.96$ on the full dataset. In other words, with MNIST, the dataset size does not significantly affect tightness. However, for CIFAR-10, smaller datasets significantly improve the audit’s tightness significantly.
Sensitivity to clipping norm.
Finally, we evaluate the impact of the hyper-parameters on the tightness of black-box audits. In the black-box setting, unlike in white-box, the adversary cannot arbitrarily insert gradients that are of the same scale as the gradient clipping norm, $C$ [31]. Thus, they are restricted to gradients that naturally arise from samples. In Figure 4, we report the $\varepsilon_{emp}$ values obtained with varying clipping norms, $C$. For computational efficiency, the results on CIFAR-10 are for models trained on datasets of size $n=1{,}000$.
@@ 5.3 Fine-Tuning the Last Layer Only
Once again, we compare the tightness of the audits with respect to the initial model parameters. In Figure 5, we report the empirical privacy leakage estimates obtained at various levels of $\varepsilon$ when the last layer is initialized to average- and worst-case (through pre-training) initial model parameters. Similar to the full model training setting, model initialization can impact audit tightness, albeit not as significantly. Specifically, at $\varepsilon=10.0$, we obtain $\varepsilon_{emp}=9.19$ and $9.33$, respectively, when the last layers are initialized to average- and worst-case initial model parameters.
Overall, the audits in this setting are considerably tight, and increasingly so for larger $\varepsilon$, as we obtain empirical privacy leakage estimates of $\varepsilon_{emp}=0.67,1.62,3.64,9.33$ for $\varepsilon=1.0,2.0,4.0,10.0$, respectively. The maximum $\varepsilon_{emp}$ across five independent runs is also much tighter, namely, $\varepsilon_{emp}=1.37,2.10,4.28,9.94$ for $\varepsilon=1.0,2.0,4.0,10.0$, respectively. Although maximum $\varepsilon_{emp}$ exceeds the theoretical $\varepsilon$ at 1.0, 2.0, and 4.0 suggests a DP violation, this is, in fact, within the standard deviation of ±$\pm 0.40,0.36,$ and $0.38$, respectively.
@@ 6 Conclusion
This paper presented a novel auditing procedure for Differentially Private Stochastic Gradient Descent (DP-SGD). By crafting worst-case initial model parameters, we achieved empirical privacy leakage estimates substantially tighter than prior work for DP-SGD in the black-box model and for natural (i.e., not adversarially crafted) datasets. At $\varepsilon=10.0$, we achieve nearly tight estimates of $\varepsilon_{emp}=7.21$ and $6.95$ for datasets consisting of 1,000 samples from MNIST and CIFAR-10, respectively. While we achieve slightly weaker estimates of $\varepsilon_{emp}=6.48$ and $4.96$ on the full datasets, these still represent a roughly 3x improvement over the estimates achieved in prior work [13, 31].
@@
Relatedly, recent work [3, 37] has begun to focus on auditing within a single training run, although it has thus far achieved only loose empirical estimates. One interesting direction would be to combine the insights from this paper (i.e., using worst-case initial model parameters) with one-shot auditing techniques. However, this may not be trivial as these techniques typically employ a large number of canary samples, each with large gradient norms, which can potentially interfere with our goal to reduce the gradient norm of “other” samples.
Furthermore, we only consider full batch gradient descent ($B=n$), i.e., without sub-sampling. Note that this is standard practice and eases auditing by enabling GDP auditing, which requires fewer models and is much less computationally intensive compared to auditing DP-SGD with sub-sampling using Privacy Loss Distributions (PLDs) [31]. Nevertheless, when auditing DP-SGD with sub-sampling in the “hidden-state” threat model where the adversary can only observe the final trained model but can insert gradient canaries, recent work has shown that there is a significant gap between the theoretical upper bound and the empirical lower bound privacy leakage achieved [10]. Even though this suggests a privacy amplification phenomenon in the hidden-state setting (which would extend to black-box as well), for general non-convex loss functions, prior work has also shown that the privacy analysis of DP-SGD with sub-sampling is tight even in the hidden-state [4]. Therefore, whether audits can be tight in the black-box threat model for DP-SGD with sub-sampling under realistic loss functions and datasets remains an open question for future work.
@@
Appendix A Model Architectures and Hyper-parameters
@@
Hyper-parameter tuning.
To achieve the best model utilities, we tune the hyper-parameters of DP-SGD (i.e., number of iterations, $T$ and learning rate, $\eta$). For MNIST, we train for $T=100$ iterations, with a learning rate of $\eta=4$ for both average-case and worst-case initial model parameters. For CIFAR-10, we train for $T=200$ iterations, with a learning rate of $\eta=2$ for the average-case initial model parameter setting. For the worst-case initial model parameter setting, we use a learning rate of $\eta=1$ instead, as this resulted in the best model performance for CIFAR-10. When fine-tuning just the last layer, we train for $T=25$ iterations, with a learning rate of $10.0$. Note that, for the experiments varying the dataset size, the learning rate was scaled accordingly so that the size of the dataset does not affect the “influence” of each sample (i.e., $\frac{\eta}{B}$). All clipping norms are set to $C=1.0$ as done in prior work [13], and batch sizes are set to the dataset size $B=n$ to ease auditing [31]. Finally, the noise multiplier $\sigma$ is calculated from the batch size and number of iterations using the Privacy loss Random Variable accountant provided by Opacus [43].